OpenStudy (anonymous):

baseballs have a radius of 9cm. find the volume of 5 balls.

OpenStudy (anonymous):

use the equation of the volume \[\frac{ 4 }{ 3 }\pi r ^{3}\] and multiply the result times 5 :D

OpenStudy (anonymous):

would the answer be 105.95?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

9^3 = 729 how the answer would be smaller than that value ?

OpenStudy (anonymous):

I don't understand what you mean

OpenStudy (anonymous):

i meant show me your steps to get 105.95

OpenStudy (anonymous):